# AI generalisation gap in comorbid sleep disorder staging

### 24 Mar 2026 v1

@@ 4.4 Clincian-in-the-loop Analysis of Sleep Transition Graphs
The Statistical Lens: Transition probability analysis revealed clear differences between stroke patients (iSLEEPS) and healthy individuals (SleepEDF-78). Stroke patients showed a higher likelihood of waking from light non-REM sleep (N2Wake: 0.629 vs. 0.469), indicating unstable sleep marked by frequent arousal rather than progression into deep restorative sleep. Clinically, this instability reduces sleep efficiency and restorative benefits, consistent with post-stroke impairment of thalamocortical and brainstem systems regulating sleep and arousal. Stroke patients showed slower N2N3 and N2REM transitions, suggesting disrupted sleep depth and REM control in the subacute stage. Healthy individuals showed stronger transitions, consistent with stable and consolidated sleep. These findings align with prior studies showing that cortical and subcortical brain lesions disrupt neural networks and lead to sleep disturbances after stroke [6].

### 26 Mar 2026 v2

@@ 4.4 Clincian-in-the-loop Analysis of Sleep Transition Graphs
The Statistical Lens: Transition probability analysis revealed clear differences between stroke patients (iSLEEPS) and healthy individuals (SleepEDF-78). Stroke patients showed a higher likelihood of waking from light non-REM sleep (N2Wake: 0.354 vs. 0.164), indicating unstable sleep marked by frequent arousal rather than progression into deep restorative sleep. Clinically, this instability reduces sleep efficiency and restorative benefits, consistent with post-stroke impairment of thalamocortical and brainstem systems regulating sleep and arousal. Stroke patients showed slower N2N3 transitions, suggesting disrupted sleep depth. Healthy individuals showed stronger transitions, consistent with stable and consolidated sleep. These findings align with prior studies showing that cortical and subcortical brain lesions disrupt neural networks and lead to sleep disturbances after stroke [6].